OpenStudy (anonymous):

SOLVING SYSTEMS USING SUBSTITUTION!!!! PLEASE HELP! What are X and Y supposed to be? I'm not good at Algebra... I don't really understand substitution. y=2x+6 3y-x+3=0

OpenStudy (anonymous):

First, you turn one of the equations into slope intercept form

OpenStudy (anonymous):

How do i do that though? Sorry, like I said I'm not good at algebra :P

OpenStudy (anonymous):

Its okay. The slope intercept form is y=mx+b y=2x+6 this equation is already in slope intercept

OpenStudy (anonymous):

next you would plug in what you got for y into the second equation 3y-x+3=0 3(2x+6)-x+3=0 and solve for x

OpenStudy (anonymous):

Oh... Okay that makes sense :)

OpenStudy (anonymous):

Urgh...sorry The 2x is supposed to be -2x and for the answer I got X=3 and Y=0. Am I right?

OpenStudy (anonymous):

x is right to get y you plug in what you got for x into the other equation

OpenStudy (anonymous):

y=2x+6 y=2(3)+6

OpenStudy (anonymous):

Okay. Thanks so much! You really helped :)

OpenStudy (anonymous):

anytime :)